The sum of an angle's complementary and supplementary is 124. Find the measure of the angle.
Answer by oberobic(2304)

a + c = 90
a + s = 180
c + s = 124
.
set up as a system of equations
1a + 1c + 0s = 90
1a + 0c + 1s = 180
0a + 1c + 1s = 124
.
add the first 2 equations and subtract the 3rd
1a + 1c + 0s = 90
1a + 0c + 1s = 180
-------------------
2a + 1c + 1s = 270
0a + 1c + 1s = 124
-----------------
2a +0c + 0s = 146
.
a = 73
.
Check this answer.
c = 90-73 = 17
s = 180-73 = 107
c+s = 124
Correct.
.
Answer: Angle is 73.
.
Done.